OpenStudy (anonymous):

What is the factored form of 3x^2-13x+10 ?

OpenStudy (anonymous):

3(x-1)(x-10/3)

OpenStudy (anonymous):

1) Find roots of 3x^2-13x+10 = 0 2) Highest term multiplier * (x-1st root) * (x-2nd root)

OpenStudy (anonymous):

3x^2-13x+10 =3x^2-10x-3x+10 ;(use middel factor theorem =3x^2-3x-10x+10 =3x(x-1)-10(x-1) =(x-1)(3x-10)